Using Kirchoff Rule find the three current I1, I2, I3 where R1= 100ohms, R2= 10ohms, R3= 5ohms, E1= 12volts and E2= 6volts


1
Expert's answer
2022-03-02T14:40:54-0500

Answer

Ohm' s law

Current

"I_1=\frac{12}{100}=0.12A"

"I_2=\frac{6}{10}=0.6A"

"I_1=\frac{6}{5}=1.2A"